OpenStudy (anonymous):

show that if the line y=kx-k touches the parabola with equation y=x^2 +5x + 10 then k=-1 or k=15

OpenStudy (anonymous):

dang that took a while

OpenStudy (anonymous):

got it though

OpenStudy (anonymous):

ok lets use the point on the graph of \[y=x^2+5x+10\] at \[(a,a^2+5a+10)\] and the slope of the tangent line is given by the derivative, namely \[y'=2x+5\] so at the point \[(a,a^2+5a+10)\] the slope is \[2a+5\] now use the point slope formula to get the equation for the line

OpenStudy (anonymous):

point is \[(a,a^2+5a+10)\] slope is \[2a+5\] equation is \[y-(a^2+5a+10)=(2a+5)(x-a)\] or \[y=(2a+5)(x-a)+a^2+5a+10\]

OpenStudy (anonymous):

thanks but im seriously confused....

OpenStudy (anonymous):

multiply out to get \[y=(2a+5)x-2a^2-5a+a^2+5a+10\] or \[y=(2a+5)x-a^2+10\] or even \[y=(2a+5)x-(a^2-10)\] and therefore \[k=2a+5=a^2-10\]

OpenStudy (anonymous):

then solve the quadratic \[2a+5=a^2-10\] \[a^2-2a-15=0\] \[(a-5)(a+3)=0\] \[x=5,a=-3\]

OpenStudy (anonymous):

if \[a=5\] then \[k=2\times 5+5=15\] and if \[a=-3\] then \[k=2\times -3+5=-1\]

OpenStudy (anonymous):

now that i have written it, ask away. i was confused at first as well

OpenStudy (anonymous):

first of all this problem is huge!!!!! and second exactly how did yo get the point to start off on?

OpenStudy (anonymous):

i mean the coordinates

OpenStudy (anonymous):

i didn't get a specific point, i just called the first coordinate "a" if the first coordinate is "a" then since we know the function, the second coordinate must be \[a^2+5a+10\]

OpenStudy (anonymous):

any point on the curve \[y=x^2+5x+10\] must look like \[(a,a^2+5a+10)\]

OpenStudy (anonymous):

oh ok thanks